how to convert inches into square foot, if i have stone marble 133 inches in length and 33 inches in breadth,and per square foot amount is Rs 135
Answer by Maths68(1474)

Conversion - Feet into Inches
=======================
1 foot = 12 inches
2 feet = 2*12 inches
3 feet = 3*12 inches........ and so on.

Result:- Multiply foot by 12 we get inches.

Conversion - Inches into feet
==============================
1 inche = 1/12 feet
2 inches = 2/12 feet
3 inches = 3/12 feet......... and so on.

Result:- Divide inches by 12 we get feet.

When we multiply feet by feet we get square feet.
When we multiply inches by inches we get square inches



Now come to your question.
It seems that you want to know total cost of stone Marbel

133 inches = 133/12 feet
33 inches = 33/12 feet
Area of the stone Marble = width * length
Area of the stone Marble = 33/12 (feet) * 133/12 (feet)
Area of the stone Marble = 4389/144 (feet)^2
Area of the stone Marble = 30.48 square feet

Cost of Stone Marble = Rs. 135 per square feet
Total Cost of stone Marble = 30.48*135 = Rs. 4114.69